OpenStudy (anonymous):

solve the following initial value problem: t(dy/dt)+5y=3t with y(1)=3

OpenStudy (jamesj):

Same thing as before. What is the integrating factor for this equation? How do you find it?

OpenStudy (anonymous):

thats the question i have on web work

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

im doing the same why which i got y=((3/7)t^7+(18/7))/t^5 but its wrong i dont know why can u help

OpenStudy (jamesj):

So what is this equation in standard form and what is its integrating factor?

OpenStudy (anonymous):

i got it the video helped me alot but i have another problem and i stock with integrate of -9t*e^-t dt can u help plz

OpenStudy (jamesj):

Following exactly the procedure of the lecture, what is the equation in standard form; i.e., the coefficient of y' is 1. It is y' + (5/t)y = 3 Now, what is the integrating factor?

OpenStudy (anonymous):

no is not the same coz my right side is -3t so when im mult by e^-t then i dont know how to integrate it ???

OpenStudy (jamesj):

I don't know where you're getting this e^(-t) business. That has got absolutely nothing to do with the solution or the method of solution.

OpenStudy (jamesj):

Is or is not your IVP this: t(dy/dt)+5y=3t with y(1)=3

OpenStudy (anonymous):

no is from different question whis is (dy/dt)-y=-9t

OpenStudy (jamesj):

Oh, your first problem.

OpenStudy (jamesj):

Ok. Yes, the integrating factor here is e^(-t)

OpenStudy (anonymous):

ok then ill mult by both side

OpenStudy (anonymous):

i got -9t*e^-t

OpenStudy (anonymous):

for right side

OpenStudy (jamesj):

so e^(-t).y' - e^(-t).y = -9t.e^(-t) i.e., [ e^(-t).y ]' = -9t.e^(-t) Now you need to integrate both sides w.r.t. t \[e^{-t}y(t) = \int\limits -9te^{-t} \ dt\] So next step is evaluating this integral.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so how to integrate the right side ?

OpenStudy (jamesj):

This is why you spent a lot of time in the calculus course before your ODE course practicing integrals. For this one, use integration by parts.

OpenStudy (anonymous):

can u show me how coz i dont remember it

OpenStudy (jamesj):

Post a new question with that integral. That kind of integration question is something a lot of people on here can help you with. I haven't time right now.

OpenStudy (anonymous):

ok thanks